Solve system by elimination method.
2x-y=12
3x+2y=-3

Answer by mananth(16949)

2x-y=12 .............1
3x+2y=-3 .............2
Eliminate y
multiply (1)by 2
Multiply (2) by 1
4x-2y=24
3x+2y= -3
Add the two equations
7x=21
/7
x=3
plug value of x in (1)
2x-y=12
6-y=12
-y=12-6
-y=6
y=-6
CHECK

2 x - y = 12
6 6 = 12
12 = 12